a lift starts falling due to its broken wire with a constant velocity of 8m/s . as safety measure the bottom has a 5m long spring on the ground and the lift stops 2 m above the ground by the spring.What is the spring constant of the spring?
1
Expert's answer
2020-02-06T08:56:00-0500

The lift falls with the constant velocity, so we can say that the initial energy of the lift is

"E_i = \frac{mv^2}{2}"

When it collides with the spring, all his kinetic energy converts to the potential energy of the spring:

"E_f = \frac{k\Delta x^2}{2}"

where "\Delta x" is the spring compression. As "E_i=E_f", we can find the spring constant:


"k = \frac{mv^2}{\Delta x^2}=\frac{8^2}{(5-2)^2}m=7.1m"